OpenStudy (anonymous):

Solve for x, where x is a real number 12x^2 = 3x could any1 please explain how to solve this step by step. I really want to be able to understand this. I already know x=0, 1/4

OpenStudy (anonymous):

subtract \(3x\) from both sides, get \[12x^2-3x=0\] divide by \(3\) get \[4x^2-x=0\]

OpenStudy (anonymous):

factor \[x(4x-1)=0\] then set each factor equal to zero and solve